OpenStudy (anonymous):

Write an equation of a line that pass through the given point and is parallel to the given line. -x + y = 12 (-3,-5)

OpenStudy (anonymous):

move the x to the right. y=x+12 since it is parallel, slopes are equal y=x+b solve for b. -5=-3+b b=-2 final equation: y=x-2